# Kishu: Time-Traveling for Computational Notebooks

### v1

@@ 3.1. Kishu Components
At the start of the notebook session, Kishu patches the session namespace in order to monitor accesses to its contents between successive cell executions (§ 4.3). It identifies the candidate Co-variables to check for updates through tracking user-referenced variable names, and passes the candidates to the Delta Detector to compute the Co-variable granularity state delta.
@@
The State Loader restores to a session state upon requested checkout. It first identifies the difference between the current session state (i.e., existing items in the namespace) and the target state according to the Checkpoint Graph, then loads only the necessary data from the Checkpoint Graph for replacing the variables that need updating (§ 5.2). If required data for checkout is missing (e.g., Kishu failed to serialize it into the Checkpoint Graph) or fails to load, the Data Restorer is invoked to restore the data.
@@ 3.2. Kishu Workflow
This section describes how users interact with Kishu during a notebook workflow. Users will attach Kishu to a notebook session upon session start; then, Kishu will monitor the namespace for state deltas to incrementally checkpoint after each cell execution (if automatic checkpointing is enabled), and perform checkout back to a previous state when requested.
@@
When initializing a notebook session, Kishu will be attached to the kernel; it will patch the namespace and initialize the Checkpoint Graph on storage.
Performing Incremental Checkpointing
After each cell execution, the Delta Detector utilizes the Patched Namespace to identify the updated Co-variables, then stores them as a new incremental checkpoint (i.e., node) in the Checkpoint Graph.
@@
Kishu will restore a previous session state on request. The State Restorer first identifies the difference between the current and target state at Co-variable granularity according to the Checkpoint Graph (i.e., which variables need to be restored), then loads only the necessary data according to the difference to complete the restoration. If necessary, the Data Restorer reconstructs data that is missing (i.e., failed to serialize during checkpointing earlier) or failed to load via fallback recomputation.
@@ 4. Accurate and Fast Delta Detection
In this section, we describe how Kishu correctly detects the Co-variable granularity state deltas necessary for incremental checkpointing and checkout (§ 2.3). We formally describe the Co-variable in § 4.1, how we correctly detect Co-variable updates in § 4.2, and how we speedup the detection process in § 4.3.
@@ 4.1. Co-variables
In Python and the Jupyter Notebook ecosystem, variables and objects are 2 distinct concepts: A variable is a named entity from which one or more objects are reachable - for example, for a list ls=[1,2,3], the list name (ls) is a variable and each of the elements (1, 2, 3) is an object. We define reachability according to references, i.e., object y is reachable from variable x if y can be accessed from x through a chain of references. Some common reachability patterns include subscripting (e.g., y = x[0]), class member (e.g., y = x.attr), and attribution (e.g., y = x.__dict__). Given our distinction between variables and objects and our definition of reachability, we now define the Co-variable as follows:
@@
A Co-variable can consist of one name (e.g., a primitive, x = 1) or multiple names from which the same object can be reached (i.e., shared references). LABEL:fig:checkpoint_variable_update shows an example—the string object ’b’ is reachable from both list ls and object obj via subscript and class member respectively, hence {ls,obj} is a Co-variable. Co-variables are self-contained by definition, i.e., there are no inter-Co-variable references. Co-variables can be modified by cell executions:
@@
For example, the Co-variable {ls,obj} in LABEL:fig:checkpoint_variable_update is modified node-wise with “ls[0] = ’e’(bottom), and is modified edge-wise with “obj.foo = ls[2](bottom-right). Co-variables can also be created (or deleted) via split and merge (right): the Co-variable {obj,ls} is deleted via a split (as obj and ls longer share references), and the Co-variable {obj, st} is created through a merge. For brevity, we collectively refer to Co-variable modifications, creations, and deletions as updates—the set of Co-variables updated by a cell execution form the execution’s state delta.
@@ 4.2. Accurate State Delta Detection
Kishu uses VarGraphs to detect Co-variable membership and updates. The VarGraph is a graph structure constructed from each variable in the namespace that captures its reachable objects. An example is shown in LABEL:fig:checkpoint_id_graph: each node in a variable’s VarGraph corresponds to a reachable object, containing the (1) object type, (2) memory address, and one of (3) pointers to other reachable objects (i.e., children) for non-primitive objects, or (4) value for primitive objects. For example, the node for the list reachable from ls contains 3 child pointers to the 3 nodes for strings ’a’, ’b’, and ’c’, and the node for string ’b’ holds its value ’b’.  [^2] The VarGraph is inspired by ElasticNotebook’s ID graph (Li et al., [n.d.]) which captures reachable objects’ memory addresses; VarGraphs uniquely contain datatypes and primitive values for additional robustness (e.g., detecting a different primitive in the same address).
@@
Co-variable membership is determined by intersecting VarGraphs. For example, in LABEL:fig:checkpoint_id_graph, ls and obj form a Co-variable as the node ’b’ is in both graphs (red).
@@ 4.3. Efficient State Delta Detection
In this section, we describe how Kishu speeds up the Co-variable update detection process. Identifying Co-variable updates across the entire global namespace via VarGraphs can be expensive (due to object traversals); hence, Kishu needs to reduce the number of Co-variables (hence the portion of the namespace) it checks after each cell execution while maintaining accuracy of detection.
@@
A Co-variable being accessed indicates the possibility of it being updated (e.g., a subscription access ls[0] = ’e’). Kishu patches the accessor, setter, and deletion methods of the global namespace (LABEL:fig:checkpoint_detect_fast) to capture variable (hence Co-variable) accesses. Kishu identifies candidate Co-variables for update checking via the captured accessed variables: if a Co-variable’s members $(x_{1},...,x_{i})$ overlaps with the accessed variables of the cell execution, then the Co-variable may have been updated (e.g., {ls,obj} in LABEL:fig:checkpoint_detect_fast) and Kishu will have to verify the update by (1) re-generating VarGraphs for its member variables, (2) comparing the VarGraphs with those before the cell execution to identify modifications, and (3) intersecting the VarGraphs amongst variables of accessed Co-variables to identify merges and splits. Otherwise, the Co-variable surely wasn’t updated and Kishu skips its check for this cell execution (e.g., {df} in LABEL:fig:checkpoint_detect_fast, greyed out). We provide a short proof of contradiction:
@@
As only a small portion of variables are accessed per cell in a typical data science notebook, Kishu significantly reduces delta detection overhead with this approach (empirically verified in § 7.6.1).
@@
The user may still use non-referencing methods such as direct C-pointer-based modifications, but these cases are rare in notebooks (found in 0 out of 60 surveyed notebooks (Li et al., [n.d.])) Some libraries, such as NumPy (Team, 2024a), do perform memory-based updates (e.g., via slicing). However, the objects are supported by Kishu as their memory-based updates are still invoked via referencing (e.g., arr[0,1] += 1, empirically verified in § 7.2.1).
@@ 5.1. Incremental Checkpointing
The Checkpoint Graph is a directed tree of (incremental) checkpoints representing the branch-based evolution of session states. It is grown (i.e., via adding nodes) with each checkpoint Kishu performs, and nodes are timestamped according to the completion time of the corresponding cell execution $t$ (which we refer as CE $t$ and node $t$ for simplicity). The Checkpoint Graph maintains a head node, which tracks the user’s current state. Each node $t$ contains the state delta consisting of (only) Co-variables updated by CE $t$. When Co-variables are stored in the Checkpoint Graph, they are versioned according to their corresponding CE $t$:
@@
Kishu creates a new node $t$ in the Checkpoint Graph after each CE $t$. The new node contains (1) the state delta of CE $t$ consisting of versioned Co-variables, (2) the code of CE $t$, and (3) the versioned Co-variables (possibly stored in previous checkpoints) accessed by CE $t$ (§ 4.3). For example, the node $t_{3}$ in LABEL:fig:restore_diff contains the code of CE $t_{3}$ (“plot=gmm.result()“) and its dependency on the versioned Co-variable from node $t_{2}$ $((gmm),t_{2})$ (dashed line), analogous to the transformation, transaction, and dependencies in database versioning. The new node $t$ is written into the Checkpoint Graph under the head node $s$, and an edge (i.e., a parent-child relationship) is added from the head node $s$ to the new node $t$ (which is now the new head node).
@@ 5.2. Efficient State Restoration
The Session State [^3] Unlike state deltas, cell code, and dependencies, session states are not stored in nodes on the Checkpoint Graph. They are inferred (by timestamp) at checkout time. at timestamp $t$ is a set of $n$ Versioned Co-variables $\{(\mathcal{X}_{1},t_{1}),...,(\mathcal{X}_{n},t_{n})\}$ such that for each $(\mathcal{X}_{j},t_{j})$, $1\leq i\leq n$:
@@
The session state at timestamp $t$ (session state $t$ for brevity) is the set of all Versioned Co-variables that are (1) available in the namespace after CE $t$ and (2) have not been overwritten by a newer Versioned Co-variable prior to CE $t$. For example, in LABEL:fig:restore_diff, the session state at timestamp $t_{3}$ (top-left) consists of the Versioned Co-variables ($(\{plot\},t_{3})$, ($(\{gmm\},t_{2})$, and ($(\{df\},t_{1})$. It does not contain ($(\{gmm\},t_{1})$ as it was overwritten by CE $t_{2}$ (gmm.fit(k=3)) which writes ($(\{gmm\},t_{2})$. Each session state $t$ dictates which Versioned Co-variables should be loaded from various nodes on the Checkpoint Graph for checkouts; to efficiently accomplish (incremental) checkout, Kishu identifies the difference between the current and target session states in terms of the (versioned) Co-variables that need to be updated. That is, some Co-variables do not need to be updated when converting the current state to the target state. They can be identified via the Checkpoint Graph:
@@
A Co-variable $\mathcal{X}$ is identical between the current state $t_{a}$ and target state $t_{b}$ if a Versioned Co-variable $(\mathcal{X},t_{x})$ exists in the session states of $t_{a}$, $t_{b}$, and $t_{c}$, where $t_{c}$ is the lowest common ancestor of node $t_{a}$ and node $t_{b}$. Otherwise, if no such $(\mathcal{X},t_{x})$ exists, then the Co-variable $\mathcal{X}$ has diverged between $t_{a}$ and $t_{b}$.
A Co-variable $\mathcal{X}$ is identical between current and target session states $t_{a}$ and $t_{b}$ if its versioned counterpart has the same version across $t_{a}$, $t_{b}$, and $t_{c}$, i.e., none of the cell executions between (1) node $t_{a}$ and node $t_{c}$ and (2) node $t_{b}$ and node $t_{c}$ updated the Co-variable $\mathcal{X}$, hence does not need to be updated when checking out from $t_{a}$ to $t_{b}$. For example, in LABEL:fig:restore_diff, if checking out from current state $t_{5}$ to target state $t_{3}$, the Co-variable {df} (blue) is identical between the states as no cell execution between (1) node $t_{1}$ and node $t_{3}$ and (2) node $t_{1}$ and node $t_{5}$ updated it.
@@ 6.1. Implementation
For seamless integration, Kishu is implemented as a separate application from the notebook process, and can be used without altering the base Jupyter application. When a session is initialized, Kishu places hooks into the kernel (pre_run_cell and post_run_cell (IPython, 2024b)) and patches the namespace (user_ns (IPython, 2024a)) (§ 3.1), allowing the standalone Kishu process to perform delta detection, write data in the namespace to storage, and overwrite data in the namespace upon checkout transparently.
@@
The Pickle protocol (i.e., __reduce__ (Foundation, 2023f)) is employed for (1) object serialization and (2) constructing VarGraphs (hence identifying Co-variables), i.e., an object y is reachable from another object x if pickle(x) includes y. As Pickle is the de-facto standard (in Python) observed by almost all data science libraries (e.g., NumPy, PyTorch (et al., 2018)), Kishu can be used for almost all use cases. Furthermore, Kishu’s per-Co-variable storage enables mixing and matching serialization libraries for coverage. Currently, Kishu will try CloudPickle (cloudpipe, 2024) first, then use Dill (Foundation, 2023g) as a fallback for Co-variables containing objects that CloudPickle fails on.
@@ 6.2. Design Considerations
Hidden and External Items
In Kishu, the session state is formally defined according to Co-variables in the user namespace (user_ns), which contains key-value pairs of variable names to their reachable objects. The session state does not include (1) local/module/hidden variables and (2) external items outside of the notebook session (e.g., on-disk files), which we do not aim to checkpoint nor restore.
@@ 7. Experimental Evaluation
    Generalizable and Robust Mechanism: Kishu is capable of identifying modifications to, and correctly restoring session states containing 146 object classes from commonly used Data Science libraries. (§ 7.2)
@@
    Low Checkpoint Storage Cost: Kishu’s incremental checkpointing optimizations result in its per-cell-execution checkpoints being up to 4.55$\times$ smaller compared to those from the next best mechanism. (§ 7.3)
@@
    Fast Incremental Checkout: Kishu’s novel incremental restoration is crucial to its sub-second checkout times — up to 8.18$\times$ and 9.02$\times$ faster than the next best mechanism for undoing cell executions and switching branches, respectively. (§ 7.5)

### v2

@@ 3.1. Kishu Components
On session start, Kishu patches the namespace to monitor accesses to its contents during cell executions (§ 4.3). It tracks user-referenced variable names to identify candidate Co-variables to check for updates in, which are passed to the Delta Detector to compute the Co-variable granularity state delta.
@@
The State Loader restores to a session state upon requested checkout. It first identifies the difference between the current session state (i.e., existing items in the namespace) and the target state according to the Checkpoint Graph, then loads only the necessary data from the Checkpoint Graph for replacing the Co-variables that need updating (§ 5.2).
@@ 3.2. Kishu Workflow
This section describes Kishu’s operations during a notebook workflow. Users will interact with Kishu using the in-Jupyter Command Palette (LABEL:fig:system_overview), such as to attach it to a started notebook session; [^3] Kishu’s interface is similar to that of ElasticNotebook; a depiction can be found in its demo paper (Li et al., 2024). However, the core functionality and techniques used are different. Kishu then monitors the namespace for state deltas to incrementally checkpoint after each cell execution (if automatic checkpointing is enabled), and perform checkout to a previous state when requested.
@@
When initializing a notebook session, init attaches Kishu to the kernel, which will patch the namespace and initialize the Checkpoint Graph on storage.
Incremental Checkpointing
After each cell execution, the Delta Detector uses the Patched Namespace to identify updated Co-variables and stores them in a new incremental checkpoint/node with a unique checkpoint_id on the Checkpoint Graph. The user may view the stored graph, checkpoints, and their IDs with log.
@@
Kishu will restore a previous session state with checkout checkpoint_id. The State Restorer identifies differing Co-variables (i.e., to restore) between the current and target states according to the Checkpoint Graph, then accordingly loads only the necessary data for restoration. If necessary, the Data Restorer reconstructs data that is missing (i.e., failed to serialize) or failed to load via fallback recomputation.
@@ 4. Accurate and Fast Delta Detection
This section describes how Kishu correctly detects Co-variable granularity state deltas necessary for incremental checkpointing and checkout (§ 2.4). We formally describe Co-variables in § 4.1 and how we detect Co-variable updates correctly (§ 4.2) efficiently (§ 4.3).
@@ 4.1. Co-variables
In Python and the Jupyter Notebook ecosystem, variables and objects are 2 distinct concepts: A variable is a named entity from which various objects are reachable—for an example list ls=[1,2,3], the list name (ls) is a variable and each list element (1, 2, 3) is an object. We define reachability reference-wise, i.e., object y is reachable from variable x if y can be accessed from x through a chain of references. Some common reachability patterns include subscripting (e.g., y = x[0]), class member (e.g., y = x.attr), and attribution (e.g., y = x.__dict__). Given our distinction between variables and objects and reachability definition, we now define Co-variables as follows:
@@
A Co-variable can consist of one name (e.g., a primitive, x=1) or multiple names from which same objects can be reached (i.e., shared references). For example, in LABEL:fig:checkpoint_variable_update, the string object ’b’ is reachable from both Pandas Series ser and object obj via subscript and class member respectively, hence {ser,obj} is a Co-variable. Co-variables are self-contained by definition, i.e., there are no inter-Co-variable references. They can be modified by cell executions:
@@
For example, the Co-variable {ser,obj} in LABEL:fig:checkpoint_variable_update is modified node-wise with an in-place update “ser.replace’’ (bottom), and edge-wise with a re-assignment “obj.foo = ser[2](bottom-right). Co-variables can also be created and deleted via split and merge (right): the Co-variable {obj,ser} is deleted via a split as obj and ser longer share references, and the Co-variable {obj, st} is created through a merge. We collectively refer to Co-variable modifications, creations, and deletions as updates; the set of Co-variables updated by a cell execution is the execution’s state delta.
@@ 4.2. Accurate State Delta Detection
Kishu uses VarGraphs to detect Co-variable membership and updates. The VarGraph is a graph structure constructed from each variable in the namespace that captures its reachable objects. An example is shown in LABEL:fig:checkpoint_id_graph: each node in a variable’s VarGraph corresponds to a reachable object, containing the (1) object type, (2) memory address, and one of (3) pointers to other reachable objects (i.e., children) for non-primitive objects, or (4) value for primitive objects. For example, the node for the list reachable from ser contains 3 child pointers to the 3 nodes for strings ’a’, ’b’, and ’c’, and the node for string ’b’ holds its value ’b’.  [^4] The VarGraph is inspired by ElasticNotebook’s ID graph (Li et al., 2023a) which captures reachable objects’ memory addresses; VarGraphs uniquely contain datatypes and primitive values for additional robustness (e.g., detecting a different primitive in the same address).
@@
Co-variable membership is determined by intersecting VarGraphs. For example, in LABEL:fig:checkpoint_id_graph, ser and obj form a Co-variable as the node ’b’ is in both graphs (red).
@@ 4.3. Efficient State Delta Detection
This section describes how Kishu speeds up Co-variable update detection. Identifying updates across the entire global namespace via VarGraphs can be expensive (due to object traversals); Kishu needs to reduce the number of Co-variables (i.e., portion of namespace) it checks after cell executions without reducing detection accuracy.
@@
Co-variable accesses indicates possible updates (e.g., via a member function call ser.replace). Kishu patches the accessor, setter, and deletion methods of the global namespace (LABEL:fig:checkpoint_detect_fast) to capture variable (hence Co-variable) accesses, which helps identify the possibly updated Co-variables: if the members of a Co-variable $(x_{1},...,x_{i})$ overlaps with the cell execution’s accessed variables, then it may have been updated (e.g., {ser,obj} in LABEL:fig:checkpoint_detect_fast): Kishu will verify the update by (1) re-generating VarGraphs for its member variables, (2) comparing the VarGraphs with those before the cell execution to identify modifications, and (3) intersecting the VarGraphs amongst variables of accessed Co-variables to identify merges and splits. Otherwise, the Co-variable surely wasn’t updated and Kishu skips its check for this cell execution (e.g., {df} in LABEL:fig:checkpoint_detect_fast, greyed out).
@@
As only a small portion of variables are accessed per cell in a typical data science notebook, Kishu significantly reduces delta detection overhead with this approach (empirically verified in § 7.6).
@@
The user may still use non-referencing methods to update data such as with direct C-pointer-based modifications, but these cases are rare in notebooks (found in 0 out of 60 surveyed notebooks (Li et al., 2023a)). Some libraries, such as NumPy (Team, 2024a), do perform memory-based updates (e.g., via slicing). However, the objects are supported by Kishu as their memory-based updates are still invoked via referencing (e.g., arr[0,1] += 1, empirically verified in § 7.2.1).
@@ 5.1. Incremental Checkpointing
The Checkpoint Graph is a directed tree of (incremental) checkpoints representing the branch-based state evolution. Nodes are added for each checkpoint Kishu performs, and timestamped according to the completion time $t$ of the corresponding cell execution (we refer to the timestamped node and cell execution as node $t$ and CE $t$, respectively, for simplicity). The Checkpoint Graph maintains a head node, which tracks the user’s current state. Each node $t$ contains the state delta consisting of (only) Co-variables updated by CE $t$. Co-variables stored in the Checkpoint Graph are versioned according to their corresponding CE $t$:
@@
Kishu creates a new node $t$ in the Checkpoint Graph after each CE $t$ containing (1) the versioned Co-variables in the state delta of CE $t$, (2) the code of CE $t$, and (3) the versioned Co-variables (possibly stored in previous checkpoints) accessed by CE $t$ (§ 4.3). For example, node $t_{3}$ in LABEL:fig:checkpoint_graph (blue) contains the code of CE $t_{3}$ (“plot=gmm.result()“) and its dependency on the versioned Co-variable from node $t_{2}$ $((gmm),t_{2})$ (dashed line). Notably, the state delta, code, and variable accesses are respectively analogous to the update, operation, and dependencies in database logging and versioning. The new node $t$ is written under the head node $s$, and and a parent-child relationship is added from the head node $s$ to the new node $t$ (now the new head node).
@@ 5.2. Efficient State Restoration
The Session State [^5] Kishu stores snapshots of Session State metadata (i.e., references to contained Co-variables) in Checkpoint Graph nodes. at timestamp $t$ is a set of $n$ Versioned Co-variables $\{(\mathcal{X}_{1},t_{1}),...,(\mathcal{X}_{n},t_{n})\}$ such that for each $(\mathcal{X}_{j},t_{j})$, $1\leq i\leq n$:
@@
The session state at timestamp $t$ (state $t$ for brevity) is the set of all Versioned Co-variables that are (1) available in the namespace after CE $t$ and (2) have not been overwritten by a newer Versioned Co-variable prior to CE $t$. For example, in LABEL:fig:restore_diff, state $t_{3}$ (top-left) consists of Versioned Co-variables ($(\{plot\},t_{3})$, ($(\{gmm\},t_{2})$, and ($(\{df\},t_{1})$. It does not contain ($(\{gmm\},t_{1})$ as it was overwritten by CE $t_{2}$ (gmm.fit(k=3)) which writes ($(\{gmm\},t_{2})$. Each state $t$ dictates which Versioned Co-variables should be loaded from various nodes on the Checkpoint Graph for checkouts; to efficiently accomplish (incremental) checkout, Kishu identifies the difference between the current and target states in terms of the (versioned) Co-variables that need to be updated: some Co-variables do not need to be updated when converting the current state to the target state. They can be identified via the Checkpoint Graph:
@@
A Co-variable $\mathcal{X}$ is identical between the current state $t_{a}$ and target state $t_{b}$ if a Versioned Co-variable $(\mathcal{X},t_{c})$ exists in the session states of $t_{a}$, $t_{b}$, and $t_{c}$, where $t_{c}$ is the lowest common ancestor of node $t_{a}$ and node $t_{b}$. Otherwise, if no such $(\mathcal{X},t_{c})$ exists, then the Co-variable $\mathcal{X}$ has diverged between $t_{a}$ and $t_{b}$.
A Co-variable $\mathcal{X}$ is identical between current and target states $t_{a}$ and $t_{b}$ if its versioned counterpart is consistent across $t_{a}$, $t_{b}$, and $t_{c}$, i.e., no cell execution between (1) node $t_{a}$ and node $t_{c}$ and (2) node $t_{b}$ and node $t_{c}$ updated $\mathcal{X}$, hence does not need updating when checking out from $t_{a}$ to $t_{b}$. For example, in LABEL:fig:restore_diff, if checking out from current state $t_{5}$ to target state $t_{3}$, the Co-variable {df} (blue) is identical between the states as no cell execution between (1) node $t_{1}$ and node $t_{3}$ and (2) node $t_{1}$ and node $t_{5}$ updated it.
@@ 6.1. Implementation
Kishu is implemented as a separate application from the notebook process, usable without altering the base Jupyter application. Upon session initialization, Kishu places hooks into the kernel (pre_run_cell and post_run_cell (IPython, 2024b)) and patches the namespace (user_ns (IPython, 2024a)) (§ 3.1), allowing the standalone Kishu process to detect state deltas, write data to storage, and overwrite data in the namespace upon checkout transparently.
@@
The Pickle protocol (i.e., __reduce__ (Foundation, 2023f)) is used for (1) object serialization and (2) constructing VarGraphs for identifying Co-variables, i.e., object y is reachable from another object x if pickle(x) includes y. As Pickle is the de-facto standard (in Python) observed by almost all libraries (e.g., NumPy, PyTorch (et al., 2018)), Kishu can be used in almost all cases. Kishu’s per-Co-variable storage also enables mixing and matching serialization libraries for coverage: Currently, Kishu will try CloudPickle (cloudpipe, 2024) first, then use Dill (Foundation, 2023g) as a fallback for Co-variables that CloudPickle fails on.
@@ 6.2. Design Considerations
Alternative Delta Detection Methods
While Kishu’s VarGraphs can generalizably detect Co-Variable updates, there are specific cases in which more efficient detection methods are applicable, for example, (1) hashing (e.g., XXH64 (xxHash, 2023)) for array-like data (Team, 2024a)) and (2) rule-based static cell (e.g., df.head) identification for skipping update detection (§ 2.2). Kishu currently uses hash-based detection for some common array-likes (e.g., TensorFlow’s tensors (TensorFlow, 2024a)), but can be extended to incorporate (1) other efficient detection methods for specific data classes and (2) rule-based cell analyses.
@@ 7. Experimental Evaluation
    Generalizable and Robust Mechanism: Kishu can identify modifications to, and correctly restore session states containing 146 object classes from common Data Science libraries. (§ 7.2)
@@
    Low Checkpoint Storage Cost: Kishu’s optimizations result in its per-cell-execution checkpoints being up to 4.55$\times$ smaller compared to those from the next best mechanism. (§ 7.3)
@@
    Fast Incremental Checkout: Kishu’s novel incremental restoration is crucial to its sub-second checkout times — up to 8.18$\times$ and 4.18$\times$ faster than the next best mechanism for undoing cell executions and switching branches, respectively. (§ 7.5)